Suppose Nancy Ford buys a brand new Chevy Cruze in September 2019. The car comes up with a standard 3 years/36,000-mile bumper-bumper warranty. Of course, Nancy will have to pay for extended warranties beyond the manufacturer’s warranty period. At the time of purchase, Bruner Motors of Stephenville offers an extended warranty at a price of $1,259. This warranty would cover years any covered parts in years 4 and 5, after the expiry of original warranty.
Explain the role of discounting in Nancy’s decision if she should or should not purchase the extended warranty.
Based on Kelly Car Reports, the estimated cost of typical repairs for Chevy Cruze is $800 in each of years 4 and 5. If Nancy uses a discount rate of 6% per year, should she purchase the extended warranty?
Would her decision be different if the discount rate were 1% per year?
Answer #1

Cost of repairs in year 4= 800 dollars

Cost of repairs in year 5= 800 dollars

Therefore we need to find the present value of both these costs, with the given discount rate of 6%.

Therefore the PV=FV/(1+R)^t

PV at year 4= 800/(1+0.06)^4= 633.67

PV at year 5= 800/(1+0.06)^5= 597.80

Therefore total= 1231.47

The extended warranty price is 1259

Therefore the extended warranty is not worth it by 28 dollars.

If the discount rate was 1 percent then;

PV at year 4= 800/(1+0.01)^4= 768.78

PV at year 5= 800/(1+0.01)^5= 761.17

Therefore total= 1529.34

At a discount rate of 1 percent, the extended warranty is worth it.

Hope this helps!